# OmniDRCA: Parallel Speech-Text Foundation Model via Dual-Resolution Speech Representations and Contrastive Alignment

### v1

@@ Abstract
Recent studies on end-to-end speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens. Existing approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis. (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation. This paper presents OmniDRCA, a parallel speech-text foundation model based on joint autoregressive modeling, featuring dual-resolution speech representations and contrastive cross-modal alignment. Our approach processes speech and text representations in parallel while enhancing audio comprehension through contrastive alignment. Experimental results on Spoken Question Answering benchmarks demonstrate that OmniDRCA establishes new state-of-the-art (SOTA) performance among parallel joint speech-text modeling based foundation models, and achieves competitive performance compared to interleaved models. Additionally, we explore the potential of extending the framework to full-duplex conversational scenarios. [^1] Our implementation will be released at https://github.com/lxchtan/OmniDRCA.
@@ 1 Introduction
Recent research efforts on E2E models to address the above challenges have focused on two primary directions: Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models. Text-Driven Speech Models refer to systems in which LLMs ingest speech representations as input, produce textual responses, and simultaneously utilize the LLM’s latent representations to synthesize audio. In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and autoregressively generating both text tokens and speech tokens. The key distinction lies in whether the generated speech tokens are fed back autoregressively into the LLM during the generation process. Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024). Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM, ensuring temporal alignment. While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations. Text-Driven Speech Models prevent LLMs from perceiving generated speech tokens during inference, thereby limiting dynamic adjustment of text generation. Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge. Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials; hence, in this work, we focus on enhancing joint speech-text modeling.
@@
In this work, we propose OmniDRCA, a novel parallel speech-text foundation model. Our architecture preserves end-to-end advantages while significantly improving speech-text semantic alignment through two key innovations: 1) Dual-Resolution Speech Representations (DRSR): DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens. For the LLM’s speech comprehension, we introduce a grouping mechanism, mapping 25Hz discrete audio tokens to 5Hz speech representations that align much better with textual granularity, thereby leveraging LLM’s capabilities more effectively and improving speech understanding. Combined speech-text embeddings form the assistant’s autoregressive input. The multimodal LLM (MLLM) predicts text tokens. For speech generation, latent representations are ungrouped and autoregressively processed by the speech refinement model (SRM) for individual speech token prediction. During inference, due to input-output resolution discrepancies, each MLLM inference triggers $k$ parallel SRM executions, where $k$ is the grouping factor. 2) Contrastive Cross-modal Alignment (CCA): Similarity between grouped speech representations and text representations is computed to optimize their mutual distances. While grouping aligns speech-text granularity, CCA further enhances semantic consistency between modalities before they enter the LLM and further enhances speech comprehension. Considering inference efficiency for varying output requirements in practical applications, we also design system prompts to control the output modes, including text-only output, parallel speech-text output, and parallel speech-text output after chain-of-modality (CoM) (Zhang et al., 2023a). Experimental results demonstrate the efficacy of CoM in improving performance. For reproducibility, we will release all source code and model checkpoints.
Our contributions are three-fold: 1) We propose OmniDRCA, a novel parallel speech-text foundation model that enhances speech-to-speech generation based on discrete speech tokens and LLMs, through dual-resolution speech representations and contrastive cross-modal alignment. 2) Comprehensive experiments demonstrate OmniDRCA’s superiority over the state-of-the-art (SOTA) parallel model LUCY (Gao et al., 2025) and competitive performance against SOTA interleaved models across multiple spoken question answering benchmarks. 3) We further explore the potential of OmniDRCA for full-duplex voice conversation. By leveraging time-division multiplexing, OmniDRCA-Duplex supports real-time responses to user interruptions with minimal performance degradation.
@@ 2 Related Work
Speech Tokenization. Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations. Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation. In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as in CosyVoice (Du et al., 2024a, b), and prioritized for linguistic content (Hsu et al., 2021). While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023). In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to its competitive semantic capabilities and support for streamlined streaming speech generation.
@@
E2E Speech Foundation Models. Existing E2E speech foundation models can be categorized into Text-Driven Speech Models and Joint Speech-Text Models. Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech. The contemporaneous Text-Driven model Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production. Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
@@
Joint Speech-Text Models explore two different architectures, Interleaved and Parallel. Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024) conduct parallel decoding. The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations. Our approach differs from Moshi with two key distinctions: (1) the $k$ tokens correspond to grouped semantic speech tokens rather than disparate acoustic tokens, and (2) our speech refinement module explicitly ingests full-sequence contextual information to enhance semantic coherence. While SLAM-Omni (Chen et al., 2024) proposes grouped semantic token prediction for speech generation, our approach designs an ungrouping mechanism and incorporates a Speech Refined Model for autoregressive generation of individual speech tokens. Another contemporaneous parallel model Kimi-Audio (KimiTeam et al., 2025) is closest to our approach. To mitigate limitations of discrete tokens in audio understanding, they introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper features, preserving both semantic and acoustic information. In contrast, our approach enhances discrete token comprehension through contrastive loss and grouping strategies, thereby improving LLM’s understanding of speech tokens while maintaining representational efficiency.
Contrastive Cross-Modal Alignment. CLIP (Radford et al., 2021) aligns image and text representations through contrastive pretraining. VILA-U (Wu et al., 2024) integrates contrastive loss and reconstruction loss during vision token pretraining to enhance feature alignment. In speech research, Ye et al. (2022) leverages multi-class N-pair contrastive loss (Sohn, 2016) for speech translation. Zheng et al. (2024) leverages InfoNCE loss (He et al., 2020) for spoken language understanding. We incorporate contrastive loss to strengthen speech-text semantic alignment. Crucially, we adopt gradient stop (Chen and He, 2021) on text embeddings during contrastive loss computation – this modification preserves speech representation learning while preventing text feature distortion.
@@ 3 Methodology
Figure 1 illustrates the architecture of OmniDRCA. The system consists of two main components: (1) a Multimodal Large Language Model (MLLM) designed to process interleaved speech-text inputs and achieve cross-modal alignment, and (2) a Speech Refined Model (SRM) specialized in high-fidelity speech synthesis. The system operates through multimodal input encoding, joint representation learning with contrastive alignment, and coordinated speech-text output generation. During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce multimodal context embeddings, and decoded into synchronized speech-text responses through SRM and the language model head.
@@ 3.1 Speech Tokenization and Detokenization
Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ HuBERT (Hsu et al., 2021) as the speech tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$, where $T$ is the speech token sequence length. For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency.
@@
Figure 1: Overview of OmniDRCA. User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text token prediction. Hidden representations are ungrouped and processed by the SRM through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor. The generated speech tokens are then converted to speech waveform by the speech detokenizer. We omit speech tokenizer and detokenizer for simplicity.
3.2 Multimodal Large Language Model (MLLM)
Parallel Joint Speech-Text Model. Leveraging the autoregressive generation capability of MLLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the MLLM at each timestep. Their embeddings are added as model input, forming a parallel speech-text architecture. Formally, the combined input embedding $c_{t}$ at timestep $t$ is computed as:
$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. The autoregressive generation process is as follows:
$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
Inner Text Streaming. Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold. We focus on performing modality alignment exclusively at the assistant end. This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
Speech Token Grouping. A crucial parameter for discrete tokens is sampling rate, which determines the input/output sequence length. GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz. Hence, our work adopts 25Hz sampling rate. To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024), a grouping mechanism is designed:
[…]
Cross-modal Contrastive Alignment.
To enable cross-modal alignment, a contrastive learning objective aligns grouped speech representations with their textual counterparts by maximizing their cosine similarity and computing a contrastive loss. This joint optimization ensures coordinated positioning of both modalities within the shared representation space.
$\mathcal{L}_{\text{cont}}=-\cos(E_{\text{text}}(\mathbf{t}),E_{\text{speech}}(\mathbf{g})),$ (4)
where $E_{\text{text}}$ and $E_{\text{speech}}$ are the mean of embeddings of text tokens and grouped speech tokens between sample pairs, respectively. Different from early applications of contrastive alignment between speech and text (Zheng et al., 2024), our approach eliminates negative sample computation and prevents mode collapse through gradient stop on the text branch.
3.3 Speech Refined Model (SRM)
SRM is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the MLLM as conditional input and incorporates contextual speech information to autoregressively generate speech tokens. While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details. To address this limitation, an ungrouping process is performed as follows: the MLLM’s final hidden state is mapped to group-sized embeddings via linear projection
$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[MLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (5)
and time splitting
$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}]\quad\text{where}\quad\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k},$ (6)
the resulting $\mathbf{H}$ serves as conditional input for SRM that autoregressively generates speech tokens. Following our practice on MLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated. SRM is trained to maximize the conditional probability:
$\mathcal{L}_{\text{SRM}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}),$ (7)
where $s_{i}$ represents the $i$-th speech token. By optimizing this objective, SRM learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from MLLM. This design enables SRM to effectively leverage the semantic and acoustic information encoded in the hidden representations of MLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
The end-to-end training objective integrates the three losses through multi-task learning paradigm:
$\mathcal{L}_{\text{total}}=\lambda\mathcal{L}_{\text{LM}}+\mu\mathcal{L}_{\text{cont}}+\gamma\mathcal{L}_{\text{SRM}},$ (8)
$\mathcal{L}_{\text{LM}}=-\sum_{i=1}^{T}\log P(t_{i}|t_{<i},\mathbf{g}),$ (9)
where $\mathcal{L}_{\text{LM}}$ is the autoregressive loss over MLLM, $\lambda$, $\mu$, and $\gamma$ are hyperparameters.
3.4 Multimodal Interaction Patterns

### v2

# DrVoice: Parallel Speech-Text Voice Conversation Model via Dual-Resolution Speech Representations

@@ Abstract
Recent studies on end-to-end speech generation with large language models (LLMs) have attracted significant community attention, with multiple works extending text-based LLMs to generate discrete speech tokens. Existing approaches primarily fall into two categories: (1) Methods that generate discrete speech tokens independently without incorporating them into the LLM’s autoregressive process, resulting in text generation being unaware of concurrent speech synthesis. (2) Models that generate interleaved or parallel speech-text tokens through joint autoregressive modeling, enabling mutual modality awareness during generation. This paper presents DrVoice, a parallel speech-text voice conversation model based on joint autoregressive modeling, featuring dual-resolution speech representations. Whereas current methods utilize mainly 12.5Hz input audio representation, our proposed dual-resolution mechanism reduces the input frequency for the LLM to 5Hz. Experimental results on Spoken Question Answering benchmarks demonstrate that DrVoice establishes new state-of-the-art (SOTA) performance among similar size speech foundation models with relative small amount of data.
@@ 1 Introduction
Recent research efforts on E2E models to address the above challenges have focused on two primary directions Chen et al. (2024): Text-Driven Speech Models (Yao et al., 2024; Li et al., 2025) and Joint Speech-Text Models. Text-Driven Speech Models refer to systems in which LLMs process speech representations as input, produce textual responses, and utilize the LLM’s representations as input to a speech decoder for speech generation. In contrast, Joint Speech-Text Models involve LLMs taking speech representations as input and generating both text tokens and speech tokens simultaneously. The key distinction lies in whether the generation of speech tokens can influence the subsequent generation of text tokens within the LLM. This feedback loop of speech tokens to LLM is present in Joint models but absent in Text-Driven models. Joint Speech-Text Models can be further categorized into interleaved speech-text modeling (Zeng et al., 2025; Zhang et al., 2024a) and parallel speech-text modeling (Défossez et al., 2024; Chen et al., 2024; KimiTeam et al., 2025). Interleaved speech-text modeling alternates speech and text representations as inputs to the LLM, while parallel speech-text modeling fuses speech and text representations before feeding them into the LLM. While both methodologies explicitly incorporate textual guidance into speech token generation to leverage the LLM’s capabilities, they have distinct limitations. The architecture of Text-Driven Speech Models creates a unidirectional information flow where text generation is completed before speech synthesis begins. This prevents the LLM from conditioning its textual output on the generated acoustic tokens, thus limiting its ability to control fine-grained paralinguistic attributes like emotion and prosody.
Joint Speech-Text Models degrade the original text generation capabilities due to speech token interference, making the preservation of text capabilities a critical challenge. Nevertheless, Joint Speech-Text Models enforce multimodal interaction and generation and empower greater potentials Chen et al. (2024); hence, in this work, we focus on enhancing joint speech-text modeling. Recently, Kimi-Audio KimiTeam et al. (2025) has set a new state-of-the-art (SOTA) in joint speech-text modeling. However, this approach faces notable challenges. It not only requires extensive data for training but also incurs significant computational costs due to its 12.5Hz audio representation. Furthermore, this high token rate creates a frequency mismatch with the much lower rate of text tokens ($\sim$3Hz) (Chen et al., 2024), which can dilute semantic information and consequently hinder the full utilization of the LLM’s core capabilities.
@@
In this work, we propose DrVoice, a novel parallel speech-text voice conversation model with Dual-Resolution Speech Representations (DRSR), which compresses the input down to 5Hz. DRSR is proposed mainly to alleviate the temporal resolution discrepancy between speech and text tokens. For speech comprehension, we introduce a grouping mechanism that maps 25Hz discrete audio tokens to 5Hz speech representations, which better match the pace of text. During generation, combined speech-text embeddings form the assistant’s autoregressive input. The hidden states captured from shared LLM layers are then passed in parallel to a text head for text token prediction, and a Speech Refined Head (SRH) to generate the corresponding ungrouped speech tokens. Due to input-output resolution discrepancies, each LLM inference triggers $k$ parallel SRH executions, where $k$ is the grouping factor. To further enhance the model’s reasoning and the coherence of its output, we incorporate a Chain-of-Modality (CoM) (Zhang et al., 2023a) strategy. This approach prompts the model to first generate a complete response in text, allowing it to structure its thoughts before engaging in parallel speech-text generation. This intermediate reasoning step serves to improve modality alignment and reduce errors. For practical applications, we design system prompts to control the output modes, enabling text-only output, direct parallel speech-text output, or the CoM-enhanced parallel output. Experimental results demonstrate the efficacy of CoM in significantly improving performance. For reproducibility, we will release all source code and model checkpoints.
Our contributions are three-fold: 1) We propose DrVoice, a novel parallel speech-text conversation model featuring Dual-Resolution Speech Representations (DRSR). This core architectural innovation effectively alleviates the temporal resolution mismatch between speech and text tokens, reducing computational costs and better preserving the LLM’s semantic processing capabilities. 2) We introduce a CoM-Mixing training strategy which acts as a curriculum, using the structured reasoning of CoM to scaffold directly speech generation task. 3) Through comprehensive experiments, we demonstrate that DrVoice achieves new state-of-the-art (SOTA) performance on multiple spoken question answering benchmarks, validating the effectiveness of our proposed architecture and strategies.
@@ 2 Related Work
Two primary directions exist for converting continuous speech signal into processable sequences: continuous representations, e.g., Whisper (Radford et al., 2022), and discrete representations. Since LLMs generate discrete tokens, continuous representations face challenges in direct integration with LLMs for speech generation. In contrast, discrete tokens enable LLMs to handle speech similar to text tokens, and are typically categorized into two categories: 1) acoustic tokens optimized for reconstructing high-quality audio through neural audio codecs (Défossez et al., 2023), and 2) semantic tokens typically derived from self-supervised pre-trained models with masked language modeling as training objective (Hassid et al., 2023) or supervised learning as in CosyVoice (Du et al., 2024a, b), and prioritized for linguistic content (Hsu et al., 2021). While acoustic tokens achieve superior acoustic fidelity in sound reconstruction, semantic tokens demonstrate stronger alignment with semantic representations (Zhang et al., 2023b), thereby facilitating more effective extension of MLLMs’ capabilities in both speech comprehension and generation (Zeng et al., 2025; Zhang et al., 2024b, 2023a; Borsos et al., 2023). In this work, we select HuBERT-derived semantic tokens (Hassid et al., 2023) over other paradigms due to their competitive semantic capabilities and compatibility with a more lightweight, diffusion-free vocoder for direct speech synthesis.
E2E Speech Foundation Models.
@@
Existing E2E speech foundation models can be categorized into Text-Driven Speech Models and Joint Speech-Text Models. Text-Driven Speech Models (Fang et al., 2024; Wang et al., 2024; Fu et al., 2025; Yao et al., 2024; Huang et al., 2025; Chen et al., 2025) integrate speech encoder, adaptor, LLM, and a streaming speech decoder, and can simultaneously generate text and speech. Qwen2.5-Omni (Xu et al., 2025) employs the Thinker-Talker architecture, with Thinker handling multimodal understanding and text generation and Talker specializing in speech token production. Since Thinker cannot receive speech tokens during generation, the framework inherently limits awareness of speech token generation states, constraining applications such as full-duplex voice conversation.
@@
Joint Speech-Text Models explore two different architectures, Interleaved and Parallel. Interleaved models (Zhang et al., 2024a; Zeng et al., 2025; Li et al., 2025; Long et al., 2025) adopt interleaved decoding to support simultaneous generation of speech and text tokens, while parallel models (Défossez et al., 2024; Xie and Wu, 2024a, b; Chen et al., 2024; KimiTeam et al., 2025) conduct parallel decoding. The interleaved model VITA-Audio Long et al. (2025) achieves efficient audio generation by introducing multiple cross-modal token prediction (MCTP) modules. It demonstrates superior performance across multiple benchmarks while maintaining high-quality speech synthesis. The parallel model Moshi (Défossez et al., 2024) employs a compact Deep Transformer to predict $k$ tokens while relying solely on current-step representations. To mitigate limitations of discrete tokens in audio understanding, Kimi-Audio (KimiTeam et al., 2025) introduce dual-tokenizer combining discrete semantic tokens with continuous Whisper Radford et al. (2023) features, preserving both semantic and acoustic information. In order to further enhance the efficiency of both training and inference and alleviate the issue of granularity misalignment between the text and speech modalities, our approach leverages DRSR to reduce the LLM input frame rate to 5Hz, without sacrificing performance.
@@ 3 Methodology
Figure 1 illustrates the architecture of DrVoice. The system consists of three main components: (1) Speech Tokenizer and Speech Encoder are designed to process the wave of speech into hidden representation, (2) a Multimodal Large Language Model (MLLM) consists of shared LLM layer, text head, and speech refined head (SRH) for token generation, and (3) Speech Detokenizer to generate wave from speech tokens. The system operates through multimodal input encoding and coordinated speech-text output generation. During inference, user inputs (text or speech) are first mapped to a unified semantic space, processed by MLLM to produce parallel speech-text responses through SRH and text head. To effectively train this system, we introduce the CoM-Mixing Training strategy.
@@ 3.1 Speech Tokenization and Detokenization
Semantic tokens have been widely used for speech tokenization (Zhang et al., 2023a; Borsos et al., 2023), due to their strong alignment with text (Zhang et al., 2023b); therefore, we employ HuBERT (Hsu et al., 2021) as the Speech Tokenizer to convert speech waveform to semantic speech token sequence $\mathbf{S}=[s_{0},s_{1},\cdots,s_{T-1}]$, where $T$ is the speech token sequence length. For speech detokenization, HiFi-GAN (Kong et al., 2020) is selected to regenerate speech waveform from speech tokens $\mathbf{S}$, since it is a non-autoregressive neural vocoder that efficiently generates high-fidelity waveforms and reduces the overall latency. Inspired by Kimi-Audio KimiTeam et al. (2025) and aiming for improved audio understanding, we utilize the Whisper-Large-v3 Speech Encoder to extract continuous audio representations. Subsequently, an Adapter is introduced to downsample the temporal resolution of these representations and align their hidden dimension with that of the LLM. Finally, the processed continuous representations are stacked with discrete ones and fed into the LLM.
@@
Figure 1: Overview of DrVoice. User speech inputs are tokenized, grouped, and encoded by the MLLM for autoregressive text and speech token prediction. The generated speech tokens are then converted to speech waveform by the speech detokenizer. Note that SRH through $k$ autoregressive forward passes to generate $k$ speech tokens, where $k$ is the grouping factor.
@@
3.2 Multimodal Large Language Model (MLLM)
Parallel Joint Speech-Text Model.
Leveraging the autoregressive generation capability of LLMs, both generated speech tokens $s_{t}$ and text tokens $t_{t}$ are iteratively fed back into the shared LLM layer at each timestep. Their embeddings are added as model input, forming a parallel speech-text architecture. Formally, the combined input embedding $c_{t}$ at timestep $t$ is computed as:
$c_{t}=E_{\text{speech}}(s_{t})+E_{\text{text}}(t_{t})$ (1)
where $E_{\text{speech}}$ and $E_{\text{text}}$ denote the embeddings of speech and text tokens, respectively. Since the lengths of speech tokens and text tokens are typically mismatched, the shorter sequence is padded with special token <|SIL|>to align them for each utterance. The autoregressive generation process is as follows:
$P(y_{t}|y_{<t},x)=\prod_{i=1}^{t}P(y_{i}|y_{<i},x)$ (2)
Inner Text Streaming.
Inspired by Moshi (Défossez et al., 2024), explicit text streaming is incorporated to aid speech generation as a common semantic scaffold. We focus on performing modality alignment exclusively at the assistant end. This design adheres to the asymmetric nature of human-machine interactions: while user inputs are typically unimodal (either text or speech), the assistant’s responses can be a coordinated multimodal output.
Speech Token Grouping.
A crucial parameter for discrete tokens is sampling rate, which determines the input/output sequence length. GLM-4-Voice (Zeng et al., 2025) investigates sampling rates from 6.25Hz to 50Hz on LibriSpeech, revealing minimal differences in Word Error Rate (WER) between 50Hz and 25Hz but significant degradation at 12.5Hz and 6.25Hz. Hence, our work adopts 25Hz sampling rate. To resolve the temporal resolution mismatch between speech signals (25Hz) and text tokenization ($\sim$3Hz) (Chen et al., 2024), a grouping mechanism is designed:
where $\mathbf{s}_{j}$ denotes speech tokens, $\|$ represents feature concatenation, and $k$ is the grouping factor determined by the ratio between speech token rates and text token rates. This design compresses the speech token length from $T$ to $T/k$ and the grouped speech representations align better with text. Different from Chen et al. (2024) that employs linear projection of audio logits into group-sized representations for parallel multi-token prediction, DrVoice strategically designs an ungrouping mechanism and incorporates a dedicated Speech Refined Head (SRH) to enable autoregressive generation of individual speech tokens.
Speech Refined Head.
SRH is proposed to enhance speech generation capabilities, which utilizes the last hidden state of the shared LLM layer (SLLM) as conditional input and incorporates contextual speech information to autoregressively generate speech tokens. While conventional speech grouping strategies–which cluster speech tokens into semantically meaningful units–have proven effective for speech recognition and understanding tasks (Chen et al., 2024; Zhang et al., 2024b), our experiments reveal their inherent limitations in generative scenarios, since speech token grouping inevitably loses some fine-grained acoustic details. To address this limitation, an ungrouping process is performed as follows: the SLLM’s final hidden state is mapped to group-sized embeddings via linear projection
$\mathbf{h}_{ug}=\mathbf{W}_{p}\mathbf{h}_{L}^{\text{[SLLM]}}\quad\text{where}\quad\mathbf{W}_{p}\in\mathbb{R}^{d_{g}\times d_{h}},$ (4)
and time splitting
$\mathbf{H}=\text{Split}_{k}(\mathbf{h}_{ug})=[\mathbf{h}_{ug}^{(1)},\mathbf{h}_{ug}^{(2)},\ldots,\mathbf{h}_{ug}^{(k)}],$ (5)
where $\mathbf{h}_{ug}^{(i)}\in\mathbb{R}^{d_{ug}/k}$, the resulting $\mathbf{H}$ serves as conditional input for SRH that autoregressively generates speech tokens. Following our practice on SLLM, representations of preceding speech tokens and $\mathbf{H}$ are aggregated. Speech token prediction is trained to maximize the conditional probability:
$\mathcal{L}_{\text{SRH}}=-\sum_{i=1}^{T}\log P(s_{i}|s_{<i},\mathbf{H}_{<i}),$ (6)
where $s_{i}$ represents the $i$-th speech token. By optimizing this objective, SRH learns to predict subsequent speech tokens conditioned on both preceding speech tokens and the rich contextual embeddings $\mathbf{H}$ derived from SLLM. This design enables SRH to effectively leverage the semantic and acoustic information encoded in the hidden representations of SLLM, thereby producing more natural and coherent speech outputs compared to conventional grouping-based approaches.
The end-to-end training objective integrates the two losses through multi-task learning paradigm:
$\mathcal{L}_{\text{MLLM}}=\lambda\mathcal{L}_{\text{TH}}+\mu\mathcal{L}_{\text{SRH}},$ (7)
$\mathcal{L}_{\text{TH}}=-\sum_{i=1}^{T}\log P(t_{i}|c_{<i},\mathbf{g}),$ (8)
where $\mathcal{L}_{\text{TH}}$ is the autoregressive loss over text head, $\lambda$ and $\mu$ are hyperparameters.
3.3 CoM-Mixing Training strategy